Each square in the grid is 0.50 mx 0.50 m in dimensions Force F2025 N and a force H-1000MFind the moments of each force about the points A.B.C and D.


1
Expert's answer
2021-08-20T12:55:56-0400





"\tau=r\cdot F\cdot \sin\alpha=d\cdot F" . So,


"\tau_{FA}=1.118\cdot2025=2264\ (N\cdot m)"


"\tau_{FB}=0.2236\cdot2025=453\ (N\cdot m)"


"\tau_{FC}=2.01246\cdot2025=4075\ (N\cdot m)"


"\tau_{FC}=1.118\cdot2025=2264\ (N\cdot m)" .


Similarly


"\tau_{HA}=949\ (N\cdot m)"


"\tau_{HB}=1897\ (N\cdot m)"


"\tau_{HC}=1265\ (N\cdot m)"


"\tau_{HD}=1739\ (N\cdot m)"